OpenStudy (anonymous):

The length of a rectangle is 4 centimeters less than twice its width. The perimeter of the rectangle is 34 cm. What are the dimensions of the rectangle

OpenStudy (xguardians):

 34 CM = (2W - 4) + (2W - 4) + W + W 34 CM = 6W - 8CM 42CM = 6W W = 7 CM Width = 7 CM Length = 2W - 4 Length = 14 - 4 Length = 10 CM Width = 7CM / Length = 10CM Hope this helps!